OpenStudy (anonymous):

help please . math Simplify and select the answer with the appropriate restrictions for the variables (x+4) (x+2) / 4x+8 a : x+2/4, x is not = -2 b: x+4/4 x is not = -2 c: 4/x+4 , x is not = -4 d: 4/x+2 , x is not = -2 6x2-3x/3x^2 2x-1/x , x is not = 0 2-3x , x is not = 0 6x-1, x is not = 0 2x-1 , x is not = 0 Simplify and select the answer with the appropriate restrictions for the variables 3x^2 +6x / x^2 -2x -8 2x^2/ x-4 , x is not = 4 3x/x-4 , x is not = to 4, -2

OpenStudy (anonymous):

(x+4)(x+2)/(4x+8) = (x+4)(x+2)/4(x+2) = (x+4)/4 pretty simple after this